OpenStudy (andriod09):

need help for two factoring questions: 1) x^3+19x+60, and 2) x^2-19x+48

OpenStudy (andriod09):

@karatechopper help me please?? and hey! long time no see!

OpenStudy (karatechopper):

hmm.. bottom one.. ( x-3) x-16)

OpenStudy (karatechopper):

Sorry I meant, (x-3)(x-16)

OpenStudy (karatechopper):

Find the factors of 48 that subtract to make 19

OpenStudy (karatechopper):

*-19

OpenStudy (andriod09):

also, how do you \[\sqrt{48}\] without a calculator?

OpenStudy (rajee_sam):

you estimate

OpenStudy (rajee_sam):

48 is closer to 49 and Sqrt of 49 is 7 so sqrt of 48 is approximately 7

OpenStudy (andriod09):

>.< i need exact answers. and i need to show work.

OpenStudy (karatechopper):

well, what are the factors of 48?

OpenStudy (karatechopper):

1,48 2,24 3,16 4,12

OpenStudy (karatechopper):

Now, which one of these factor pairs have a number that has a square root? 3,16 and 4,12

OpenStudy (karatechopper):

Lets try 4,12 the square root of 4 is two. So that goes outside the radical. Making the radical become.. \[2\sqrt{12}\] but we notice that the 12 can still be reduced. So we know that that factor pair does not work.

OpenStudy (karatechopper):

Lets try 3,16 Whats the square root of 16? >>4. So 4 would be outside the radical..making it look like this: \[4\sqrt{3}\]

OpenStudy (karatechopper):

we notice that that radical can't be reduced more... THEREFORE.. \[\sqrt{48}=4\sqrt{3}\]